# Bayesian Perspective for Orientation Estimation in Cryo-EM and Cryo-ET

### v1

@@ Acknowledgment
Consider $h$ a real-values function defined on a closed and bounded set, $\Theta$, of $\mathbb{R}^{p}$. If there exists a unique solution $\theta^{*}$ satisfying
$\theta^{*}=\underset{\theta\in\Theta}{\operatorname*{arg\,max}}\ h(\theta),$ (A.2)
@@
[…]
Let us define the function $h(g)$ as follows:
$h(g)\triangleq-\left\|y-\Pi(g^{-1}\cdot\overline{V})\right\|^{2}.$ (A.4)
Under the assumption of the proposition that $\hat{g}_{\mathsf{MAP}}$ is unique, and given the definition of $\hat{g}_{\mathsf{MAP}}$ in (2.10), the function $h$ satisfies the conditions of Theorem A.1, and its unique maximizer is $\hat{g}_{\mathsf{MAP}}$, i.e.,
$\hat{g}_{\mathsf{MAP}}=\underset{g\in\mathsf{SO}(3)}{\operatorname*{arg\,max}}\ h(g).$ (A.5)
Then, applying Theorem A.1 for $\lambda=1/2\sigma^{2}$ and using (A.5), we obtain:
$\lim_{\sigma\to 0}\hat{g}_{\mathrm{MMSE}}$ $=\lim_{\sigma\to 0}\frac{\int_{\mathsf{SO}(3)}g\exp\Big{(}-\left\|y-\Pi(g^{-1}\cdot\overline{V})\right\|^{2}/2\sigma^{2}\Big{)}\mathrm{d}g}{\int_{\mathsf{SO}(3)}\exp\Big{(}-\left\|y-\Pi(g^{-1}\cdot\overline{V})\right\|^{2}/2\sigma^{2}\Big{)}\mathrm{d}g}$
$=\lim_{\lambda\to\infty}\frac{\int_{\mathsf{SO}(3)}g\exp\Big{(}-\lambda\left\|y-\Pi(g^{-1}\cdot\overline{V})\right\|^{2}\Big{)}\mathrm{d}g}{\int_{\mathsf{SO}(3)}\exp\Big{(}-\lambda\left\|y-\Pi(g^{-1}\cdot\overline{V})\right\|^{2}\Big{)}\mathrm{d}g}$
$=\underset{g\in\mathsf{SO}(3)}{\operatorname*{arg\,max}}\ h(g)=\hat{g}_{\mathsf{MAP}},$ (A.6)
@@
Appendix B Isotropic Gaussian distribution of SO(3) rotations
$f(\omega)=\frac{1-\cos\omega}{\pi}\sum_{\ell=0}^{\infty}(2\ell+1)e^{-\ell(\ell+1)\eta^{2}}\frac{\sin\left((\ell+1/2)\omega\right)}{\sin(\omega/2)}.$ (B.1)
Notably, the uniform distribution on $\mathsf{SO}(3)$, denoted $\mathcal{U}_{\mathsf{SO}(3)}$, corresponds to uniformly sampled axes and is characterized by the density function $f(\omega)=\frac{1-\cos\omega}{\pi}$, which represents the limiting case of $\eta\to\infty$ in (B.1).
$f(\omega)\simeq\frac{1-\cos\omega}{\sqrt{\pi}}\eta^{-\frac{3}{2}}e^{\left(\frac{\eta}{4}-\frac{\left(\omega/2\right)^{2}}{\eta}\right)}\left(\frac{\omega-e^{-\pi^{2}/\eta}\left(\left(\omega-2\pi\right)e^{\pi\omega/\eta}+\left(\omega+2\pi\right)e^{-\pi\omega/\eta}\right)}{2\sin\left(\frac{\omega}{2}\right)}\right).$ (B.2)
Appendix C Proof of Proposition 4.1
[…]
where
$p_{i}^{(t)}(g)=\frac{\exp\Big{(}-\left\|y_{i}-\left(g^{-1}\cdot\hat{V}^{(t)}\right)\right\|^{2}/2\sigma^{2}\Big{)}}{\int_{g\in\mathsf{SO}(3)}\exp\Big{(}-\left\|y_{i}-\left(g^{-1}\cdot\hat{V}^{(t)}\right)\right\|^{2}/2\sigma^{2}\Big{)}\mathrm{d}g}.$ (C.2)
As the action by the group element $g$ preserves the norm, we have,
$\left\|y_{i}-\left(g^{-1}\cdot V\right)\right\|^{2}=\left\|g\cdot{y_{i}}-V\right\|^{2}=\left\|g\cdot y_{i}\right\|^{2}+\left\|V\right\|^{2}-2{\,\langle{g\cdot y_{i}},V\rangle}.$ (C.3)
Thus, substituting (C.3) into (C.1), leads to:
[…]
As the term $\left\|g\cdot y_{i}\right\|^{2}$ is independent of $V$, (C.4) can be simplified to:
[…]
Let us denote by $h(V)$ the right hand side of (C.5):
$h(V)\triangleq\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)\left\{\left\|V\right\|^{2}-2{\,\langle{g\cdot y_{i}},V\rangle}\right\}\mathrm{d}g.$ (C.6)
Taking the first-order condition with respect to $V$ gives
$\frac{\partial}{\partial V}h(V)=2\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)\left(V-g\cdot y_{i}\right)\mathrm{d}g=0.$ (C.7)
As $\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)=1$ for every $i$, we can simplify (C.7) to obtain:
$V=\frac{1}{M}\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}gp_{i}^{(t)}(g)\mathrm{d}g.$ (C.8)
The integral in the right-hand-side of (C.8) is nothing but $\hat{g}_{\mathsf{MMSE},i,t}$ as defined in (4.5). Thus, we obtain
$\hat{V}^{(t+1)}=\frac{1}{M}\sum_{i=0}^{M-1}\hat{g}_{\mathsf{MMSE},i,t}\cdot y_{i},$ (C.9)
which proves the proposition.

### v2

@@ Appendix B Alignment between two noisy volumes
Consider $h$ a real-valued function defined on a closed and bounded set $\Theta$, of $\mathbb{R}^{p}$. Let $\pi$ be a positive density on $\Theta$. If there exists a unique solution $\theta^{*}$ satisfying
$\theta^{*}=\underset{\theta\in\Theta}{\operatorname*{arg\,max}}\ h(\theta),$ (D.3)
@@
$\underset{\lambda\to\infty}{\lim}\frac{\int_{\Theta}\theta e^{\lambda h(\theta)}\pi(\theta)\mathrm{d}\theta}{\int_{\Theta}e^{\lambda h(\theta)}\pi(\theta)\mathrm{d}\theta}=\theta^{*},$ (D.4)
provided $h$ is continuous at $\theta^{*}$.
Let $h(g)=-\left\|y-\Pi(g\circ{V})\right\|^{2}$ and $\pi(g)\mathrm{d}g=\mathrm{d}\Lambda(g)$. We immediately get
$\lim_{\sigma\to 0}\hat{g}_{\mathrm{relax}}$ $=\lim_{\sigma\to 0}\frac{\int_{\mathsf{SO}(3)}g\exp\Big(-\left\|y-\Pi(g\circ{V})\right\|^{2}/2\sigma^{2}\Big)\mathrm{d}\Lambda\left(g\right)}{\int_{\mathsf{SO}(3)}\exp\Big(-\left\|y-\Pi(g\circ{V})\right\|^{2}/2\sigma^{2}\Big)\mathrm{d}\Lambda\left(g\right)}$
$=\lim_{\lambda\to\infty}\frac{\int_{\mathsf{SO}(3)}g\exp\Big(-\lambda\left\|y-\Pi(g\circ{V})\right\|^{2}\Big)\mathrm{d}\Lambda(g)}{\int_{\mathsf{SO}(3)}\exp\Big(-\lambda\left\|y-\Pi(g\circ{V})\right\|^{2}\Big)\mathrm{d}\Lambda(g)}$
$=\underset{g\in\mathsf{SO}(3)}{\operatorname*{arg\,max}}\ h(g)=\hat{g}_{\mathsf{MLE}}.$ (D.5)
Finally, we note that $\hat{g}_{\mathsf{MMSE}}$ is obtained by applying the Orthogonal Procrustes procedure to project $\hat{g}_{\mathsf{relax}}$ back onto $\mathsf{SO}(3)$. Since $\hat{g}_{\mathsf{MLE}}$ is itself an element of $\mathsf{SO}(3)$, it follows naturally that
$\lim_{\sigma\to 0}\hat{g}_{\mathsf{MMSE}}=\hat{g}_{\mathsf{MLE}}.$ (D.6)
Appendix E Isotropic Gaussian distribution of SO(3) rotations
$f(\omega)=\frac{1-\cos\omega}{\pi}\sum_{\ell=0}^{\infty}(2\ell+1)e^{-\ell(\ell+1)\eta^{2}}\frac{\sin\left((\ell+1/2)\omega\right)}{\sin(\omega/2)}.$ (E.1)
Notably, the uniform distribution on $\mathsf{SO}(3)$, denoted $\mathcal{U}_{\mathsf{SO}(3)}$, corresponds to uniformly sampled axes and is characterized by the density function $f(\omega)=\frac{1-\cos\omega}{\pi}$, which represents the limiting case of $\eta\to\infty$ in (E.1).
$f(\omega)\simeq\frac{1-\cos\omega}{\sqrt{\pi}}\eta^{-\frac{3}{2}}e^{\left(\frac{\eta}{4}-\frac{\left(\omega/2\right)^{2}}{\eta}\right)}\left(\frac{\omega-e^{-\pi^{2}/\eta}\left(\left(\omega-2\pi\right)e^{\pi\omega/\eta}+\left(\omega+2\pi\right)e^{-\pi\omega/\eta}\right)}{2\sin\left(\frac{\omega}{2}\right)}\right).$ (E.2)
Appendix F Proof of Proposition 4.1
$\hat{V}^{(t+1)}=\underset{V}{\operatorname*{arg\,min}}\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)\left\|y_{i}-\left(g\circ V\right)\right\|^{2}\mathrm{d}g,$ (F.1)
where
$p_{i}^{(t)}(g)=\frac{\exp\Big(-\left\|y_{i}-\left(g\circ\hat{V}^{(t)}\right)\right\|^{2}/2\sigma^{2}\Big)}{\int_{g\in\mathsf{SO}(3)}\exp\Big(-\left\|y_{i}-\left(g\circ\hat{V}^{(t)}\right)\right\|^{2}/2\sigma^{2}\Big)\mathrm{d}g}.$ (F.2)
As the action by the group element $g$ preserves the norm, we have,
$\left\|y_{i}-\left(g\circ V\right)\right\|^{2}=\left\|g^{-1}\circ{y_{i}}-V\right\|^{2}=\left\|g^{-1}\circ y_{i}\right\|^{2}+\left\|V\right\|^{2}-2{\,\langle{g^{-1}\circ y_{i}},V\rangle}.$ (F.3)
Thus, substituting (F.3) into (F.1), leads to:
$\hat{V}^{(t+1)}=\underset{V}{\operatorname*{arg\,min}}\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)\left\{\left\|g^{-1}\circ y_{i}\right\|^{2}+\left\|V\right\|^{2}-2{\,\langle{g^{-1}\circ y_{i}},V\rangle}\right\}\mathrm{d}g.$ (F.4)
As the term $\left\|g\circ y_{i}\right\|^{2}$ is independent of $V$, (F.4) can be simplified to:
$\hat{V}^{(t+1)}=\underset{V}{\operatorname*{arg\,min}}\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)\left\{\left\|V\right\|^{2}-2{\,\langle{g^{-1}\circ y_{i}},V\rangle}\right\}\mathrm{d}g.$ (F.5)
Let us denote by $h(V)$ the right hand side of (F.5):
$h(V)\triangleq\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)\left\{\left\|V\right\|^{2}-2{\,\langle{g^{-1}\circ y_{i}},V\rangle}\right\}\mathrm{d}g.$ (F.6)
Taking the first-order condition with respect to $V$ gives
$\frac{\partial}{\partial V}h(V)=2\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)\left(V-g^{-1}\circ y_{i}\right)\mathrm{d}g=0.$ (F.7)
As $\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)=1$ for every $i$, we can simplify (F.7) to obtain:
$V=$ $\frac{1}{M}\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)\cdot(g^{-1}\circ y_{i})$
$=$ $\frac{1}{M}\sum_{i=0}^{M-1}\big(\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)\cdot g^{-1}\mathrm{d}g\big)\circ y_{i}.$ (F.8)
Thus, using the definition (4.5) we obtain
$\hat{V}^{(t+1)}=\frac{1}{M}\sum_{i=0}^{M-1}\hat{\mathfrak{g}}_{\mathsf{MMSE},i,t}\circ y_{i},$ (F.9)